OpenStudy (chris215):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=x-7/x+3 and g(x)= -3x-7/x-1

OpenStudy (chris215):

f(g(x))= -3(x-7/x-3) / (x-7/x+3)-1 is this correct?

OpenStudy (anonymous):

in f(g(x) put gx in place of x in f(x)

OpenStudy (chris215):

I already did that part so now do I have to plug in fx for gx?

OpenStudy (anonymous):

yes.u need to find f(g(x) as well as g(f(x)

OpenStudy (chris215):

\[-3\frac { \frac{ x-7 }{ x-3} }{ \frac{ x-7 }{ x+3 }-1}\]

OpenStudy (chris215):

is this correct? but with -3 in the numerator

OpenStudy (anonymous):

nope u need to write fx in gx only at place of x

OpenStudy (chris215):

I already did that part

Nnesha (nnesha):

3 is not at the bottom \[\huge\rm f(x) = \frac{ \color{red}{x}-7 }{ \color{red}{x}+3 }\] replace x by \[\large\rm \color{red}{ \frac{ -3x-7 }{ \color{Red}{x-1} }}\]

Nnesha (nnesha):

which one is it g(f(x)) or f(g(x) ?

Nnesha (nnesha):

nvm i got it's g(f(x))

Nnesha (nnesha):

you*

Nnesha (nnesha):

http://prntscr.com/778ajm

Nnesha (nnesha):

\[\huge\rm \frac { -3(\frac{ x-7 }{ x-3})-7 }{ \frac{ x-7 }{ x+3 }-1}\] that's how it suppose to be

OpenStudy (chris215):

oh ok thanks so much

Nnesha (nnesha):

uhh now you have to solve this (who made that q ) :(

OpenStudy (chris215):

yeah how should I start solving it?

OpenStudy (chris215):

should I distrubute the -3?

Nnesha (nnesha):

gimme a sec let me find shortcut ;D

OpenStudy (chris215):

ok thanks :)

Nnesha (nnesha):

yeah you have to distribute by -3 -.^

OpenStudy (chris215):

\[\frac{ \frac{ -2(5x-21) }{ x-3 } }{ \frac{ x-7 }{ x+3 }-1}\]

OpenStudy (chris215):

idk if I did it right lol

Nnesha (nnesha):

-2 ? where is it come from ?

OpenStudy (chris215):

ugh i typed it wrong

OpenStudy (chris215):

\[-\frac{ 2(5x - 21 }{ x-3 }\]

Nnesha (nnesha):

\[\huge\rm \frac { -3(\frac{ x-7 }{ x-3})-7 }{ \frac{ x-7 }{ x\color{Red}{-}3 }-1}\] it's negative 3

Nnesha (nnesha):

 multiply (x-7) by -3